# Canonical decompositions of 3-connected graphs

### v2

@@ 1.6. Background on 2-separations
Every edge $t_{1}t_{2}$ of $T$, when directed from $t_{1}$ to $t_{2}$ say, induces the separation $(X_{1},X_{2})$ of $G$ for $X_{i}:=\bigcup_{t\in T_{i}}V_{t}$, where $T_{i}$ is the component of $T-t_{1}t_{2}$ that contains $t_{i}$, provided that both $X_{1}\setminus X_{2}$ and $X_{2}\setminus X_{1}$ are non-empty. We call these separations the induced separations of $(T,{{\mathcal{V}}})$. In this paper, all tree-decompositions have the property that all their edges induce separations. The separator of $(X_{1},X_{2})$ is the adhesion set $V_{t_{1}}\cap V_{t_{2}}$, which is why we also refer to the adhesion sets of $(T,{{\mathcal{V}}})$ as the separators of $(T,{{\mathcal{V}}})$.
@@
We shall use the (2-Separation Theorem). of Tutte [63] with the total-nestedness description by Cunningham and Edmonds [20], which we recall below with the notation most suitable here. Let us say that a 2-separation of a graph is totally nested if it is nested with every 2-separation of the graph.
Theorem 1.6.1 (2-Separation Theorem).
@@
    If the torso associated with a $t\in T$ is 3-connected or a cycle, then the adhesion sets induced by the edges $st\in E(T)$ are pairwise distinct.
We refer to the tree-decompositions provided by the (2-Separation Theorem). as Tutte-decompositions as customary. We provide a proof of the (2-Separation Theorem). in Section 3.3. A far reaching extension of the 2-separation theorem (that also applies to infinite matroids and extends [27, 54]) was proved by Aigner-Horev, Diestel and Postle [1].
1.7. Apex-decompositions
@@
Let $G$ be a graph and $v\in V(G)$ a vertex. An apex-decomposition of $G$ with centre $v$ is a star-decomposition ${\mathcal{A}}$ of $G-v$ of adhesion two such that its central torso is a cycle $O$ and all adhesion sets are pairwise distinct. We refer to $O$ as the central torso-cycle of ${\mathcal{A}}$. The intersection of a leaf-bag $B_{\ell}$ of ${\mathcal{A}}$ with the centre-bag of ${\mathcal{A}}$ is the adhesion set of $B_{\ell}$. We call the edges of $O$ that are spanned by adhesion sets of leaf-bags bold, and all other edges of $O$ are timid. Note that the timid edges of $O$ exist in $G$, while possibly some but not necessarily all bold edges of $O$ exist in $G$. An apex-decomposition is 2-connected if all its leaf-bags are 2-connected.
@@
Figure 1.7. In the depicted situation, the separator of the pseudo-reduction induced by $\ell$ consists of the red edges
@@
Each leaf $\ell$ of $S$ induces the 2-separation $(B_{\ell},\bigcup_{t\in S-\ell}B_{t})=:(X_{\ell},Y_{\ell})$ of $G-v$. The pseudo-reduction of $(X_{\ell},Y_{\ell})$ is the mixed 3-separation $(X,Y)$ of $G$ defined as follows (see Figure 1.7):
@@
A set $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}$ of mixed-separations of $G$ is a star with leaves $A_{i}$ if $(A_{i},B_{i})\leqslant(B_{j},A_{j})$ for all distinct indices $i,j\in I$. In this context, we also refer to $A_{i}$ as the leaf-side of $(A_{i},B_{i})$.
@@
The set of pseudo-reductions of the separations induced by ${\mathcal{A}}$ is a star of mixed 3-separations of $G$, which we call the tri-star of ${\mathcal{A}}$. We will show in Section 1.7 that the elements of the tri-star are tri-separations, provided that $O$ essentially is not too short.
@@
Let $\sigma$ be a star of mixed-separations of a graph $G$. We say that a mixed-separation $(A,B)$ of $G$ interlaces $\sigma$ if for every $(C,D)\in\sigma$ either $(C,D)<(A,B)$ or $(C,D)<(B,A)$; see Figure 1.8.
Figure 1.8. A mixed-separation $(A,B)$ interlaces a star $\{\,(C_{i},D_{i}):i\in[4]\,\}$
@@
Let us consider the 2-connected graph $G^{\prime}:=G-v$. Let ${{\mathcal{T}}}=(T,{{\mathcal{V}}})$ be the Tutte-decomposition of $G$ provided by the (2-Separation Theorem). (1.6.1). Since $(A,B)$ and $(C,D)$ cross in $G$ so that the centre consists of $v$ and all links have size one, their induced mixed 2-separations of $G^{\prime}$ cross with empty centre and all links of size one. Hence there is a bag $V_{t}\in{{\mathcal{V}}}$ whose torso is a cycle $O$ that alternates between $S(A,B)-v$ and $S(C,D)-v$.
@@
It remains to show that both $(A,B)$ and $(C,D)$ interlace the tri-star of ${\mathcal{A}}$. By symmetry, it suffices to show this for $(A,B)$. Consider any pseudo-reduction $(X,Y)$ induced by a leaf $\ell$ of $S$. Without loss of generality, the leaf-bag $B_{\ell}$ is included in $A$. We claim that $(X,Y)\leqslant(A,B)$. The side $X$ is obtained from $B_{\ell}$ by possibly adding the vertex $v$. Since $v$ lies in the separator of $(A,B)$ by assumption, this gives $X\subseteq A$. The side $Y$ is obtained from $(V(G)\setminus B_{\ell})\cup\{a_{1},a_{2}\}$ by possibly removing some of the vertices in the adhesion set $\{a_{1},a_{2}\}$ of $B_{\ell}$. Since $B$ is included in $(V(G)\setminus B_{\ell})\cup\{a_{1},a_{2}\}$, it suffices to show that $a_{i}\notin Y$ implies $a_{i}\notin B$ for both $i=1,2$. If $a_{i}$ is not contained in $Y$, then this is because $a_{i}$ is edgy, i.e. $a_{i}$ has just one neighbour outside $B_{\ell}$ in $G$. If $B$ contains $a_{i}$, then $a_{i}$ has two neighbours in $B$ since $(A,B)$ is a tri-separation. The neighbour of $a_{i}$ in $B\cap B_{\ell}$ can only be $a_{3-i}$ since $B\cap B_{\ell}\subseteq\{a_{1},a_{2}\}$. But then $O$ cannot alternate between the separators of $(A,B)$ and $(C,D)$, as $\{a_{1},a_{2}\}\subseteq S(A,B)$ but $a_{1}a_{2}$ is a bold edge of $O$ and therefore cannot lie in $S(C,D)$. ∎
We label the edges of the central torso-cycle $O$ of an apex-decomposition ${\mathcal{A}}$ with the letters b or t, depending on whether they are bold or timid, respectively. The cyclic sequence of these letters is the type of $O$. When $O$ has type btbt, we say that $O$ has the type ${\tt btbt}^{-}$ if $O$ additionally has a timid edge both of whose endvertices are not adjacent to $v$; otherwise we say that $O$ has the type ${\tt btbt}^{+}$.
@@
When we say that we assume Section 1.7 with crossing tri-separations, this means that we also assume that the tri-star of ${\mathcal{A}}$ is interlaced by two crossing tri-separations of $G$ such that their separators intersect exactly in the vertex $v$ and such that $O$ alternates between the two separators (minus $v$).
@@
In Section 1.7 with crossing tri-separations, $O$ does not have type bbt, bbb or ${\tt btbt}^{-}$.
@@
Case bbt. Since $O$ alternates between $S(A,B)-v$ and $S(C,D)-v$ by assumption, but neither $S(A,B)$ nor $S(C,D)$ can contain a bold edge of $O$, it follows that one of $S(A,B)$ and $S(C,D)$ contains both endvertices of the timid edge of $O$, say $S(A,B)$ contains them. But then one of $A\setminus B$ or $B\setminus A$ is empty, contradicting that $(A,B)$ is a tri-separation.
@@
Assume Section 1.7 with crossing tri-separations. Then the tri-star of ${\mathcal{A}}$ consists of totally-nested strong nontrivial tri-separations.
We prove Section 1.7 across the next two sections.
1.8. Proof of Section 1.7: Special cases
Figure 1.9. The situation in the proof of Section 1.8
@@
Assume Section 1.7 with crossing tri-separations. If $O$ has length three, then the tri-star of ${\mathcal{A}}$ consists of totally-nested strong nontrivial tri-separations, $O$ has type ttt or btt, and $v$ is adjacent to all vertices of $O$.
@@
Let $(A,B)$ and $(C,D)$ denote the two crossing tri-separations from the assumptions. If $O$ has type ttt, then the tri-star of ${\mathcal{A}}$ is empty, and we are done. So $O$ has at least one bold edge. By Section 1.7, $O$ does not have type bbt or bbb, so $O$ must have type btt. Let $x_{1}x_{2}x_{3}:=O$ so that $x_{1}x_{2}$ is the bold edge of $O$; see Figure 1.9.
Since $O$ alternates between $S(A,B)-v$ and $S(C,D)-v$, since $x_{1}x_{2}$ cannot lie in $S(A,B)$ or $S(C,D)$, and since neither separator can contain both endvertices of $x_{2}x_{3}$ or of $x_{3}x_{1}$, we find that $x_{3}$ cannot lie in either separator. Thus only four elements of $V(O)\sqcup E(O)$ can lie in $S(A,B)$ or $S(C,D)$. Hence the separators of $(A,B)$ and $(C,D)$ are determined up to symmetry, say $S(A,B)-v=\{x_{2},x_{1}x_{3}\}$ and $S(C,D)-v=\{x_{1},x_{2}x_{3}\}$; see Figure 1.9. Using that $(A,B)$ and $(C,D)$ are tri-separations, we infer that $v$ is adjacent to all three vertices $x_{1},x_{2},x_{3}$ of $O$. We also recall that $v$ has a neighbour in the leaf-bag of ${\mathcal{A}}$ other than $x_{1}$ and $x_{2}$ by Section 1.7. Hence $(E,F):=(V(G)-x_{3},\{x_{1},x_{2},x_{3},v\})$ with $S(E,F)=\{v,x_{1},x_{2}\}$ is a strong tri-separation of $G$ and the unique element of the tri-star of ${\mathcal{A}}$. The tri-separation $(E,F)$ is nontrivial by Section 1.2. Since $G[F\setminus E]$ is a $K_{1}$, the tri-separation $(E,F)$ is half-connected.
We claim that $(E,F)$ is totally nested. By Section 1.5, it suffices to show that $(E,F)$ is externally tri-connected. As $v$ is joined to $x_{1}$ and $x_{2}$ by edges, it remains to show that $S(E,F)$ is externally tri-connected around $v$. For this, note that $x_{1}$ and $x_{2}$ are joined by three internally disjoint paths avoiding $v$: we find two paths in the 2-connected leaf-bag, and the third path is $x_{1}x_{3}x_{2}$. ∎
@@
Assume Section 1.7. If $O$ has type ${\tt btbt}^{+}$, then the tri-star of ${\mathcal{A}}$ consists of totally-nested strong nontrivial tri-separations.
@@
Since $O$ has type btbt and ${\mathcal{A}}$ is 2-connected, $G-v$ is obtained from the disjoint union of two 2-connected graphs $X$ and $Y$ by adding a matching of size two between $X$ and $Y$. Call the two matching edges $e_{1}$ and $e_{2}$. We denote the endvertices of $e_{i}$ in $X$ and in $Y$ by $x_{i}$ and $y_{i}$, respectively, for both $i=1,2$. Since the torso-cycle $O$ has type ${\tt btbt}^{+}$, we can use Section 1.7 to find that $O$ has two opposite vertices that are adjacent to $v$ or else the two ends of some bold edge of $O$ are neighbours of $v$ while no endvertex of the other bold edge is adjacent to $v$. We consider the two cases separately.
Case 1: $O$ has two opposite vertices that are adjacent to $v$ in $G$. Without loss of generality, $x_{1}$ and $y_{2}$ are adjacent to $v$ in $G$. By symmetry, it suffices to show that the pseudo-reduction induced by $X$ (viewing $X$ as a leaf-bag of ${\mathcal{A}}$) is a totally-nested nontrivial tri-separation of $G$. Since $x_{1}v$ is an edge in $G$, and since $v$ has a neighbour in $X\setminus\{x_{1},x_{2}\}$ by Section 1.7, the pseudo-reduction induced by $X$ is either $(X+v,Y\cup\{v,x_{1},x_{2}\})$ with separator $\{x_{1},v,x_{2}\}$ or $(X+v,Y\cup\{v,x_{1}\})$ with separator $\{x_{1},v,e_{2}\}$, depending on whether $x_{2}v$ is an edge in $G$ or not, respectively. In either case, since $v$ also has a neighbour in $Y\setminus\{y_{1},y_{2}\}$ by Section 1.7, we have a half-connected nontrivial tri-separation that is strong. So by Section 1.5, it remains to show external tri-connectivity.
@@
Subcase 1b: $x_{2}v$ is not an edge in $G$. Then the separator is $\{x_{1},v,e_{2}\}$. The endvertex $x_{2}$ of $e_{2}$ is $v$-free as $x_{2}v$ is not an edge. For external tri-connectivity around $v$, we find two internally disjoint $x_{1}$$x_{2}$ paths in $X$. The endvertex $y_{2}$ of $e_{2}$ is $x_{1}$-free as $x_{1}y_{2}$ is not an edge in $G$. For external tri-connectivity around $x_{1}$, we find two internally disjoint $y_{2}$$v$ paths in $G-x_{1}-e_{2}$, one through $Y$ and via a neighbour of $v$ in $Y\setminus\{y_{1},y_{2}\}$ (which exists by Section 1.7), and the second one is $y_{2}v$.
Case 2: $x_{1}$ and $x_{2}$ are adjacent to $v$ while $y_{1}$ and $y_{2}$ are not, say. First, we consider the pseudo-reduction induced by $X$. This is $(X+v,Y\cup\{v,x_{1},x_{2}\})$ with separator $\{v,x_{1},x_{2}\}$. We verify external tri-connectivity as in Subcase 1a. The pseudo-reduction induced by $Y$ is $(X+v,Y+v)$ or $(X+v,Y)$, depending on whether $v$ has more than one or just one neighbour in $Y\setminus\{y_{1},y_{2}\}$, respectively. In either case, we have a strong half-connected nontrivial tri-separation, and by Section 1.5 it remains to verify external tri-connectivity. For $(X+v,Y+v)$ we notice that $y_{1}$ and $y_{2}$ are $v$-free and find two internally disjoint $y_{1}$$y_{2}$ paths in $Y$, which suffices as $v$ is the only vertex in the separator $\{v,e_{1},e_{2}\}$. For $(X+v,Y)$ there is nothing to show as its separator consists of three edges. ∎
1.9. Proof of Section 1.7: General case
In the previous section, we have seen that the conclusion of Section 1.7 holds if $O$ has length three (Section 1.8) or if $O$ has type btbt (Section 1.7 with Section 1.8). In this section, we show that the conclusion of Section 1.7 also holds if $O$ has length at least four and $O$ does not have type btbt, thereby completing the proof of Section 1.7.

### v3

@@ 7 Background on 2-separations
Recall that in this paper, every separation $(A,B)$ is required to satisfy $A\setminus B\neq\emptyset\neq B\setminus A$. Every edge $t_{1}t_{2}$ of $T$, when directed from $t_{1}$ to $t_{2}$ say, induces the separation $(X_{1},X_{2})$ of $G$ for $X_{i}:=\bigcup_{t\in T_{i}}V_{t}$, where $T_{i}$ is the component of $T-t_{1}t_{2}$ that contains $t_{i}$, provided that both $X_{1}\setminus X_{2}$ and $X_{2}\setminus X_{1}$ are non-empty. We call these separations the induced separations of $(T,{{\mathcal{V}}})$. In this paper, all tree-decompositions have the property that all their edges induce separations. The separator of $(X_{1},X_{2})$ is the adhesion set $V_{t_{1}}\cap V_{t_{2}}$, which is why we also refer to the adhesion sets of $(T,{{\mathcal{V}}})$ as the separators of $(T,{{\mathcal{V}}})$.
@@
We shall use the (2-Separation Theorem). of Tutte [66] with the total-nestedness description by Cunningham and Edmonds [21], which we recall below with the notation most suitable here. Let us say that a 2-separation of a graph is totally nested if it is nested with every 2-separation of the graph. A cycle $O$ alternates between two sets $X,Y\subseteq V(G)\cup E(G)$ if $O$ contains exactly two elements $x_{1},x_{2}$ from $X$ and exactly two elements $y_{1},y_{2}$ from $Y$ and $O$ induces the cycling ordering $x_{1}<y_{1}<x_{2}<y_{2}$ on these four elements (possibly after swapping the names of the $x_{i}$ or of the $y_{i}$). We emphasise that $O$ may have arbitrary length and still alternate.
Theorem 7.1 (2-Separation Theorem).
@@
    If the torso associated with a node $t\in T$ is 3-connected or a cycle, then the adhesion sets induced by the edges $st\in E(T)$ are pairwise distinct.
We refer to the unique tree-decomposition of $G$ provided by the (2-Separation Theorem). as the Tutte-decomposition of $G$ as customary. We provide a proof of the (2-Separation Theorem). in Section 1. A far reaching extension of the 2-separation theorem (that also applies to infinite matroids and extends [28, 57]) was proved by Aigner-Horev, Diestel and Postle [1].
8 Apex-decompositions
@@
Let $G$ be a graph and $v\in V(G)$ a vertex. An apex-decomposition of $G$ with centre $v$ is a star-decomposition ${\mathcal{A}}$ of $G-v$ of adhesion two such that its central torso is a cycle $O$ and all adhesion sets are pairwise distinct. We refer to $O$ as the central torso-cycle of ${\mathcal{A}}$. The intersection of a leaf-bag $B_{\ell}$ of ${\mathcal{A}}$ with the centre-bag of ${\mathcal{A}}$ is the adhesion set of $B_{\ell}$. We call the edges of $O$ that are spanned by adhesion sets of leaf-bags bold, and all other edges of $O$ are timid. Note that the timid edges of $O$ exist in $G$, while possibly some but not necessarily all bold edges of $O$ exist in $G$. An apex-decomposition is 2-connected if all its leaf-bags $G[B_{\ell}]$ are 2-connected. We stress that we really consider the leaf-bags $G[B_{\ell}]$ rather than the torsos, and that this will be important in the proofs to come.
@@
Figure 10: In the depicted situation, the separator of the pseudo-reduction induced by $\ell$ consists of the red edges
@@
Each leaf $\ell$ of $S$ induces the 2-separation $(B_{\ell},\bigcup_{t\in S-\ell}B_{t})=:(X_{\ell},Y_{\ell})$ of $G-v$. The pseudo-reduction of $(X_{\ell},Y_{\ell})$ is the mixed 3-separation $(X,Y)$ of $G$ defined as follows (see Figure 10):
@@
A set $\sigma=\{\,(A_{i},B_{i}):i\in I\,\}$ of mixed-separations of $G$ is a star with leaves $A_{i}$ if $(A_{i},B_{i})<(B_{j},A_{j})$ for all distinct indices $i,j\in I$. In this context, we also refer to $A_{i}$ as the leaf-side of $(A_{i},B_{i})$.
@@
The set of pseudo-reductions of the separations induced by ${\mathcal{A}}$ is a star of mixed 3-separations of $G$, which we call the tri-star of ${\mathcal{A}}$. We will show in Section 8 that the elements of the tri-star are tri-separations, provided that $O$ essentially is not too short.
@@
Let $\sigma$ be a star of mixed-separations of a graph $G$. We say that a mixed-separation $(A,B)$ of $G$ interlaces $\sigma$ if for every $(C,D)\in\sigma$ either $(C,D)<(A,B)$ or $(C,D)<(B,A)$; see Figure 11. Note that for every $(C,D)\in\sigma$, neither $(C,D)$ nor $(D,C)$ interlaces $\sigma$.
Figure 11: A mixed-separation $(A,B)$ interlaces a star $\{\,(C_{i},D_{i}):i\in[4]\,\}$
@@
Let us consider the 2-connected graph $G^{\prime}:=G-v$. Let ${{\mathcal{T}}}=(T,{{\mathcal{V}}})$ be the Tutte-decomposition of $G$ provided by the (2-Separation Theorem). (7.1). Since $(A,B)$ and $(C,D)$ cross in $G$ so that the centre consists of $v$ and all links have size one, their induced mixed 2-separations of $G^{\prime}$ cross with empty centre and all links of size one. Hence there is a bag $V_{t}\in{{\mathcal{V}}}$ whose torso is a cycle $O$ that alternates between $S(A,B)-v$ and $S(C,D)-v$.
@@
It remains to show that both $(A,B)$ and $(C,D)$ interlace the tri-star of ${\mathcal{A}}$. By symmetry, it suffices to show this for $(A,B)$. Consider any pseudo-reduction $(X,Y)$ induced by a leaf $\ell$ of $S$.
Claim 8.0.1.
The adhesion-set of $B_{\ell}$ is distinct from the set of endvertices of every edge in $S(A,B)\cup S(C,D)$.
Proof of Claim. Otherwise, since $G[B_{\ell}]$ is 2-connected, it would provide a path joining the endvertices of the edge while avoiding the separator that contains the edge, a contradiction. $\diamondsuit$
Hence, without loss of generality, the leaf-bag $B_{\ell}$ is included in $A$. We claim that $(X,Y)\leqslant(A,B)$. The side $X$ is obtained from $B_{\ell}$ by possibly adding the vertex $v$. Since $v$ lies in the separator of $(A,B)$ by assumption, this gives $X\subseteq A$. The side $Y$ is obtained from $(V(G)\setminus B_{\ell})\cup\{a_{1},a_{2}\}$ by possibly removing some of the vertices in the adhesion set $\{a_{1},a_{2}\}$ of $B_{\ell}$. Since $B$ is included in $(V(G)\setminus B_{\ell})\cup\{a_{1},a_{2}\}$, it suffices to show that $a_{i}\notin Y$ implies $a_{i}\notin B$ for both $i=1,2$. If $a_{i}$ is not contained in $Y$, then this is because $a_{i}$ is edgy, i.e. $a_{i}$ has just one neighbour outside $B_{\ell}$ in $G$. If $B$ contains $a_{i}$, then $a_{i}$ has two neighbours in $B$ since $(A,B)$ is a tri-separation, and these two neighbours lie in $G-v$ as $a_{i}$, being edgy, is not adjacent to $v$. The neighbour of $a_{i}$ in $B\cap B_{\ell}$ can only be $a_{3-i}$ since $B\cap B_{\ell}\subseteq\{a_{1},a_{2}\}$. But then $O$ cannot alternate between the separators of $(A,B)$ and $(C,D)$, as $\{a_{1},a_{2}\}\subseteq S(A,B)$ but $a_{1}a_{2}$ is a bold edge of $O$ and therefore cannot lie in $S(C,D)$. ∎
We label the edges of the central torso-cycle $O$ of an apex-decomposition ${\mathcal{A}}$ with the letters b or t, depending on whether they are bold or timid, respectively. The cyclic sequence of these letters is the type of $O$. For example, $O$ has type btbt if $O$ is a 4-cycle $O=v_{0}v_{1}v_{2}v_{3}v_{0}$ and there is an index $i$ such that the edges along either $v_{i}v_{i+1}v_{i+2}v_{i+3}v_{i+4}$ or $v_{i}v_{i-1}v_{i-2}v_{i-3}v_{i-4}$ (using indices in $\mathbb{Z}_{4}$) read btbt. When $O$ has type btbt, we say that $O$ has the type ${\tt btbt}^{-}$ if $O$ additionally has a timid edge both of whose endvertices are not adjacent to $v$; otherwise we say that $O$ has the type ${\tt btbt}^{+}$.
@@
When we say that we assume Section 8 with crossing tri-separations, this means that we also assume that the tri-star of ${\mathcal{A}}$ is interlaced by two crossing tri-separations of $G$ such that their separators intersect exactly in the vertex $v$ and such that $O$ alternates between the two separators (minus $v$).
@@
In Section 8 with crossing tri-separations $(A,B)$ and $(C,D)$, bold edges of $O$ cannot belong to $S(A,B)\cup S(C,D)$.
Proof.
The same as for Claim 8.0.1. ∎
[…]
In Section 8 with crossing tri-separations, $O$ does not have type bbt, bbb or ${\tt btbt}^{-}$.
@@
Case bbt. Since $O$ alternates between $S(A,B)-v$ and $S(C,D)-v$ by assumption, but neither $S(A,B)$ nor $S(C,D)$ can contain a bold edge of $O$ by Section 8, it follows that one of $S(A,B)$ and $S(C,D)$ contains both endvertices of the timid edge of $O$, say $S(A,B)$ contains them. But then one of $A\setminus B$ or $B\setminus A$ is empty, contradicting that $(A,B)$ is a tri-separation.
@@
Assume Section 8 with crossing tri-separations. Then the tri-star of ${\mathcal{A}}$ consists of totally-nested strong nontrivial tri-separations.
We prove Section 8 across the next two sections.
9 Proof of Section 8: Special cases
Figure 12: The situation in the proof of Section 9
@@ 9 Proof of Section 8: Special cases
Assume Section 8 with crossing tri-separations. If $O$ has length three, then the tri-star of ${\mathcal{A}}$ consists of totally-nested strong nontrivial tri-separations, $O$ has type ttt or btt, and $v$ is adjacent to all vertices of $O$.
@@
Let $(A,B)$ and $(C,D)$ denote the two crossing tri-separations from the assumptions. If $O$ has type ttt, then the tri-star of ${\mathcal{A}}$ is empty, and we are done. So $O$ has at least one bold edge. By Section 8, $O$ does not have type bbt or bbb, so $O$ must have type btt. Let $x_{1}x_{2}x_{3}:=O$ so that $x_{1}x_{2}$ is the bold edge of $O$; see Figure 12.
Since $O$ alternates between $S(A,B)-v$ and $S(C,D)-v$, since $x_{1}x_{2}$ cannot lie in $S(A,B)$ or $S(C,D)$, and since neither separator can contain both endvertices of $x_{2}x_{3}$ or of $x_{3}x_{1}$, we find that $x_{3}$ cannot lie in either separator. Thus only four elements of $V(O)\sqcup E(O)$ can lie in $S(A,B)$ or $S(C,D)$. Hence the separators of $(A,B)$ and $(C,D)$ are determined up to symmetry, say $S(A,B)-v=\{x_{2},x_{1}x_{3}\}$ and $S(C,D)-v=\{x_{1},x_{2}x_{3}\}$; see Figure 12. Using that $(A,B)$ and $(C,D)$ are tri-separations, we infer that $v$ is adjacent to all three vertices $x_{1},x_{2},x_{3}$ of $O$. We also recall that $v$ has a neighbour in the leaf-bag of ${\mathcal{A}}$ other than $x_{1}$ and $x_{2}$ by Section 8. Hence $(E,F):=(V(G)-x_{3},\{x_{1},x_{2},x_{3},v\})$ with $S(E,F)=\{v,x_{1},x_{2}\}$ is a strong tri-separation of $G$ and the unique element of the tri-star of ${\mathcal{A}}$. It follows immediately from the definition of $(E,F)$ that $(E,F)$ is nontrivial. Since $G[F\setminus E]$ is a $K_{1}$, the tri-separation $(E,F)$ is half-connected.
We claim that $(E,F)$ is totally nested. By Section 6, it suffices to show that $(E,F)$ is externally tri-connected. As $v$ is joined to $x_{1}$ and $x_{2}$ by edges, it remains to show that $S(E,F)$ is externally tri-connected around $v$. For this, note that $x_{1}$ and $x_{2}$ are joined by three internally disjoint paths avoiding $v$: we find two paths in the 2-connected leaf-bag, and the third path is $x_{1}x_{3}x_{2}$. ∎
@@
Assume Section 8. If $O$ has type ${\tt btbt}^{+}$, then the tri-star of ${\mathcal{A}}$ consists of totally-nested strong nontrivial tri-separations.
@@
Since $O$ has type btbt and ${\mathcal{A}}$ is 2-connected, $G-v$ is obtained from the disjoint union of two 2-connected graphs $X$ and $Y$ by adding a matching of size two between $X$ and $Y$. Call the two matching edges $e_{1}$ and $e_{2}$. We denote the endvertices of $e_{i}$ in $X$ and in $Y$ by $x_{i}$ and $y_{i}$, respectively, for both $i=1,2$. Since the torso-cycle $O$ has type ${\tt btbt}^{+}$, we can use Section 8 to find that $O$ has two opposite vertices that are adjacent to $v$ or else the two ends of some bold edge of $O$ are neighbours of $v$ while no endvertex of the other bold edge is adjacent to $v$. We consider the two cases separately.
Case 1: $O$ has two opposite vertices that are adjacent to $v$ in $G$. Without loss of generality, $x_{1}$ and $y_{2}$ are adjacent to $v$ in $G$. By symmetry, it suffices to show that the pseudo-reduction induced by $X$ (viewing $X$ as a leaf-bag of ${\mathcal{A}}$) is a totally-nested nontrivial tri-separation of $G$. Since $x_{1}v$ is an edge in $G$, and since $v$ has a neighbour in $X\setminus\{x_{1},x_{2}\}$ by Section 8, the pseudo-reduction induced by $X$ is either $(X+v,Y\cup\{v,x_{1},x_{2}\})$ with separator $\{x_{1},v,x_{2}\}$ or $(X+v,Y\cup\{v,x_{1}\})$ with separator $\{x_{1},v,e_{2}\}$, depending on whether $x_{2}v$ is an edge in $G$ or not, respectively. In either case, since $v$ also has a neighbour in $Y\setminus\{y_{1},y_{2}\}$ by Section 8, we have a half-connected nontrivial tri-separation that is strong. So by Section 6, it remains to show external tri-connectivity.
@@
Subcase 1b: $x_{2}v$ is not an edge in $G$. Then the separator is $\{x_{1},v,e_{2}\}$. The endvertex $x_{2}$ of $e_{2}$ is $v$-free as $x_{2}v$ is not an edge. For external tri-connectivity around $v$, we find two internally disjoint $x_{1}$$x_{2}$ paths in $X$. The endvertex $y_{2}$ of $e_{2}$ is $x_{1}$-free as $x_{1}y_{2}$ is not an edge in $G$. For external tri-connectivity around $x_{1}$, we find two internally disjoint $y_{2}$$v$ paths in $G-x_{1}-e_{2}$, one through $Y$ and via a neighbour of $v$ in $Y\setminus\{y_{1},y_{2}\}$ (which exists by Section 8), and the second one is $y_{2}v$.
Case 2: $x_{1}$ and $x_{2}$ are adjacent to $v$ while $y_{1}$ and $y_{2}$ are not, say. First, we consider the pseudo-reduction induced by $X$. This is $(X+v,Y\cup\{v,x_{1},x_{2}\})$ with separator $\{v,x_{1},x_{2}\}$. We verify external tri-connectivity as in Subcase 1a. The pseudo-reduction induced by $Y$ is $(X+v,Y+v)$ or $(X+v,Y)$, depending on whether $v$ has more than one or just one neighbour in $Y\setminus\{y_{1},y_{2}\}$, respectively. In either case, we have a strong half-connected nontrivial tri-separation, and by Section 6 it remains to verify external tri-connectivity. For $(X+v,Y+v)$ we notice that $y_{1}$ and $y_{2}$ are $v$-free and find two internally disjoint $y_{1}$$y_{2}$ paths in $Y$, which suffices as $v$ is the only vertex in the separator $\{v,e_{1},e_{2}\}$. For $(X+v,Y)$ there is nothing to show as its separator consists of three edges. ∎
10 Proof of Section 8: General case
In the previous section, we have seen that the conclusion of Section 8 holds if $O$ has length three (Section 9) or if $O$ has type btbt (Section 8 with Section 9). In this section, we show that the conclusion of Section 8 also holds if $O$ has length at least four and $O$ does not have type btbt, thereby completing the proof of Section 8.